# XXLTraffic: Expanding and Extremely Long Traffic Dataset for Ultra-Dynamic Forecasting Challenges

### v1

@@ Abstract
Traffic forecasting is crucial for smart cities and intelligent transportation initiatives, where deep learning has made significant progress in modeling complex spatio-temporal patterns in recent years. However, current public datasets have limitations in reflecting the ultra-dynamic nature of real-world scenarios, characterized by continuously evolving infrastructures, varying temporal distributions, and temporal gaps due to sensor downtimes or changes in traffic patterns. These limitations inevitably restrict the practical applicability of existing traffic forecasting datasets. To bridge this gap, we present XXLTraffic, the largest available public traffic dataset with the longest timespan and increasing number of sensor nodes over the multiple years observed in the data, curated to support research in ultra-dynamic forecasting. Our benchmark includes both typical time-series forecasting settings with hourly and daily aggregated data and novel configurations that introduce gaps and down-sample the training size to better simulate practical constraints. We anticipate the new XXLTraffic will provide a fresh perspective for the time-series and traffic forecasting communities. It would also offer a robust platform for developing and evaluating models designed to tackle ultra-dynamic and extremely long forecasting problems. Our dataset supplements existing spatio-temporal data resources and leads to new research directions in this domain.
1 Introduction
Traffic prediction plays a vital role in intelligent transportation systems by offering accurate and dependable forecasts of future traffic conditions that serve as critical references for route design, urban construction, and road management. The long-term spatio-temporal relationship is essential and irreplaceable for traffic forecasting, i.e., the long-term relationship can not only provide context that helps understand anomalies or irregularities in short-term data. but also help identify patterns and trends over different timescales, such as daily, weekly, monthly, and yearly cycles. In recent years, significant work has focused on short-term and long-term traffic flow prediction. Various deep learning techniques, such as GNNs, emphasize extracting spatial relationships in traffic networks [13], while Transformer variants and upgraded architectures and methods focus on capturing short-term and long-term temporal relationships [29]. Existing work has achieved promising performance in both short-term and long-term predictions, but this is limited to relatively simple experimental setups.
The inevitable tendency in intelligent transportation systems is to design models that incorporate more realistic circumstances rather than just simple experimental conditions. This requires the use of data sets that accurately reflect real-world scenarios. This serves as a strong incentive for us to create and develop XXLTraffic, i.e., why does XXLTraffic focus on ultra-dynamic forecasting? Why is it necessary to include extremely long-term data? We will introduce the commonalities and developments of existing datasets, as well as the difficulties in establishing XXLTraffic more specifically.
@@ 1 Introduction
Figure 1: Our dataset are evolving and longer than existing datasets. Existing datasets are either limited by short temporal spans or insufficient spatial nodes. In contrast, our dataset features an evolving growth of spatial nodes and spans over 20 years.
@@
Recent Advances in Expanding Traffic Datasets. Real-world traffic scenarios necessitate more complex prediction settings, involving extended temporal horizons or broader spatial coverage in experiments. In the temporal domain, new settings are typically proposed based on previously published work rather than introducing new datasets:  [30] and [11] expanded input and output lengths up to four times on existing datasets.From a spatial perspective, [4] published a dataset with nodes growing annually and provided an evolving network to support new node predictions. [33] proposed a continual learning framework with pattern expansion mechanisms based on [4]. Additionally, [27] and [25] offered larger-scale spatial node datasets to support subsequent researchers. Recent work has explored longer temporal step experimental settings and released traffic datasets spanning up to five years and thousands of nodes. However, in specific scenarios, such as future traffic prediction for highway planning, these data and experimental settings fall short. As shown in Figure 1, most existing datasets have limitations in temporal span, which inspired us to develop a dataset for expanding and extremely long traffic forecasting. Highway construction typically spans several years, requiring predictions several years in advance. These extremely long prediction scenarios necessitate traffic datasets with even longer temporal spans for support. As the temporal span extends, urban infrastructure development and road construction can lead to shifts in traffic patterns, resulting in an evolving domain shift. This observation motivated us to provide an expanding and extremely long traffic dataset. Additionally, the combination of these factors enables the extraction of more temporal patterns from extremely long sequences, allowing for the possibility of longer input sequences.
@@
Challenges. The ultra-dynamic challenge encompasses three key aspects: (1) Continuously evolving states of the underlying spatio-temporal infrastructures, characterized by an expanding number of nodes over the years. This continuous growth introduces complexity as the infrastructure adapts and expands. (2) Evolving temporal distributions over an extremely long observation period, which is crucial for extremely long forecasting. This requires models to adapt to changes in patterns and trends over extensive temporal spans. (3) Ultra spatio-temporal dynamicity with temporal gaps. These gaps can occur due to various reasons such as sensor data becoming unavailable during road construction, maintenance activities, or shifts in data distributions. This dynamic nature necessitates robust modeling techniques to handle inconsistencies and interruptions in data.
@@
Contribution. We have constructed a traffic dataset with an exceptionally long temporal span and broader regional coverage, providing aggregated data and benchmarking, as well as a benchmarking setup considering extremely long prediction scenarios for future exploration:
@@
  • We present an experimental setup with temporal gaps for extremely long prediction with gaps scenarios and provide a benchmark of aggregated versions of hourly and daily datasets.
  • We provide the exploration of input features through evolving temporal distributions over an extremely long observation period. Additionally, our datasets support zero-shot forecasting for new sensors.
2 Preliminaries
@@
Traffic data primarily consists of vehicle flow detection data collected by sensors distributed across various locations in the traffic network. It is generally represented by $X_{i}$in $\mathbb{R}^{N\times T\times C}$, where $T$ denotes the time steps, $N$ denotes the number of sensors, and $C$ denotes the number of features.
@@
$[X_{t-(\alpha-1)},...,X_{t-1},X_{t}]\rightarrow[X_{t+1},X_{t+2},...,X_{t+\beta}],$ (1)
@@
This task mainly focuses on long-sequence time series prediction, which includes the traffic dataset. As shown Table 1, the sequence length can reach up to 2880 steps.
@@
Based on Equation 1, the observation and prediction are not adjacent but are instead separated by a gap period $g$, as shown in the following formula.
$[X_{t-(\alpha-1)},...,X_{t-1},X_{t}]\rightarrow[X_{t+g+1},X_{t+g+2},...,X_{t+g+\beta}],$ (2)
3 Gaps and Comparison with Existing Traffic Datasets
Figure 2: Our expanding and extremely long prediction addresses the existing limitations in both short-term and long-term predictions.
As shown in Table 1, existing traffic prediction work can easily be divided into short-term and long-term settings. The short-term setting originated from the STGCN[41] work, while the long-term setting was first introduced by LSTNet [14] and subsequently established as a widely adopted experimental framework by Informer [45]. In recent years, short-term prediction typically has a maximum step length of 12 steps, while long-term prediction reaches up to 720 steps. However, works such as Witran [11] and DAN [19] recognized the need for even longer step predictions in practical applications, extending the length to a maximum of four times the typical length. Despite the differences in step lengths, their observed and predicted values are concatenated tightly together, as shown in Equation 1. To accommodate complex real-world scenarios, such as highway route planning predictions, it is necessary to introduce a gap of several years between observation and prediction. Typically, existing datasets lack the temporal coverage required to support gaps exceeding one year.At the same time, predicting several years in advance also implies the need to forecast traffic for sensors at new locations, taking into account the evolving nature of the road network. Even when such coverage is available, works like [33] and [4] utilize evolving datasets but do not provide sufficient data to train models for extended durations. To overcome these gaps, our Expanding and Extremely Long Traffic Dataset robustly supports these complex scenarios.
Table 1: Summary of recent short-term traffic forecasting and long-term multivariate forecasting
@@ 3 Gaps and Comparison with Existing Traffic Datasets
Short-term STGCN [41] {3,6,9,12}
DCRNN [20], GWN [40], BTF [3], DMSTGCN [10], {3,6,12}
GTS [28], STGODE [7], PM-MemNet [16], STAEFormer [23]
AGCRN [1], STSGCN [32], ,DSTAGNN [15] D2STGNN [31], {12}
DyHSL [44],PDFormer [12], MultiSPANS [47], GMSDR [22]
Long-term MTGNN [39],LSTNet [14] {3,6,12,24}
ARU [6] {12,24,48,168,336}
LogSparse_Trans [18] {24,48,72,96,120,144,168,192}
AST [38] {8,24,168,336}
SSDNet [21] {20,24,30,138}
Informer [45], Autoformer [37], FEDformer [46], {24,48,96,192,336,720}
Linear [17], Triformer [5], Pyraformer [24]
DSformer [42],DeepTime [35],DLinear [43]
Witran [11] {168, 336, 720, 1440, 2880}
DAN [19] {288, 672, 1440}
4 The XXLTraffic Datasets
4.1 Data Collection
We obtained the expanding and extremely long traffic sensor data from the California Department of Transportation (CalTrans) Performance Measurement System [^1] https://pems.dot.ca.gov/ (PeMS) [2]. PeMS is an online platform that collects traffic data from 19,561 sensors distributed across California state highways. These sensor locations are divided into nine districts. We downloaded all the raw data for these nine districts from the initial data release up to March 20, 2024. The system automatically generates a daily data file for each district, containing data from all sensors within each district. We have stored the complete raw data files in an open-source repository for quick access, which will be released after the publication.
Table 2: Comparison between our XXLTraffic dataset and the existing traffic datasets.
@@ 4.1 Data Collection
DCRNN [20] METR-LA 34,272 207 5 mins 4 months 03/2012 - 06/2012
@@
LSTNet [14] Traffic 17,544 862 1 hour 2 years 01/2015 - 12/2016
STGCN [41] PEMSD7(M) 12,672 228 5 mins 2 months 05/2012 - 06/2012
@@
ASTGCN [9] PEMSD4-I 17,002 228 5 mins 2 months 01/2018 - 02/2018
@@
STSGCN [32] PEMS03 26,208 358 5 mins 11 months 01/2018 - 11/2018
@@
Large-ST [25] CA 525,888 8,600 5 mins 5 years 01/2017 - 12/2021
@@
As illustrated in Table 2, our collected dataset significantly exceeds existing datasets in terms of both temporal coverage and the number of spatial nodes. The dataset sample will be available on: https://github.com/cruiseresearchgroup/XXLTraffic, which includes the raw data, sensor metadata (containing sensor IDs, geographical coordinates, associated road information, etc.), the data processing pipeline code, and the processed datasets.
4.2 Data Preprocessing
Based on the 23 years of raw data we collected, we conducted rigorous data filtering and aggregation. The PeMS system has continuously evolved, expanding from a few sensors in 2001 to over 4,000 sensors in some districts today. To support our setting of extremely long forecasting with gaps, we selected a subset of sensors that were installed in the early stages and have consistently collected new data up to the present(named gap dataset), which is shown in the Appendix. This extensive gap dataset effectively underpins the extremely long forecasting with gaps demonstrated in Figure 2. Utilizing the gap dataset, we performed both hourly and daily aggregations, which will be employed for gap-free long-term forecasting benchmarking. We will provide standard long-term forecasting benchmarks for both the hourly and daily datasets.
4.3 Data Overview
The XXLTraffic dataset is distributed across highways in the state of California, as illustrated in the Figure 3(a). The nine colors represent nine districts. From Figures 3(b),  3(c), and 3(d), we can clearly observe the evolutionary growth of the sensors. The sensors are extensively distributed across both urban and suburban areas, offering diverse modalities. Additionally, the sensors are densely interconnected, enabling the formation of a high-quality traffic graph dataset.

### v2

# XXLTraffic: Expanding and Extremely Long Traffic forecasting beyond test adaptation

@@ Abstract.
Traffic forecasting is crucial for smart cities and intelligent transportation initiatives, where deep learning has made significant progress in modeling complex spatio-temporal patterns in recent years. However, current public datasets have limitations in reflecting the distribution shift nature of real-world scenarios, characterized by continuously evolving infrastructures, varying temporal distributions, and long temporal gaps due to sensor downtimes or changes in traffic patterns. These limitations inevitably restrict the practical applicability of existing traffic forecasting datasets. To bridge this gap, we present XXLTraffic, largest available public traffic dataset with the longest timespan collected from Los Angeles, USA, and New South Wales, Australia, curated to support research in extremely long forecasting beyond test adaptation. Our benchmark includes both typical time-series forecasting settings with hourly and daily aggregated data and novel configurations that introduce gaps and down-sample the training size to better simulate practical constraints. We anticipate the new XXLTraffic will provide a fresh perspective for the time-series and traffic forecasting communities. It would also offer a robust platform for developing and evaluating models designed to tackle the extremely long forecasting problems beyond test adaptation. Our dataset supplements existing spatio-temporal data resources and leads to new research directions in this domain.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06copyright: acmlicensed
1. Introduction
Rapid global population growth and vehicle proliferation have intensified urban traffic congestion. As cities expand and personal transportation reliance grows, strain on road networks leads to longer commutes, higher fuel consumption, and increased emissions. Accurate traffic prediction is vital for intelligent transportation systems, informing strategies to mitigate congestion and enhance mobility through improved route planning and urban development. Effective forecasting requires capturing long-term spatio-temporal relationships in traffic data. Long-term analysis provides context for anomalies in short-term patterns and reveals trends influenced by population cycles, seasonal shifts, and yearly vehicle usage changes. These insights are crucial for developing robust models that adapt to evolving urban traffic dynamics due to demographic and vehicular changes.
In recent years, significant work has focused on both short-term and long-term traffic flow prediction. Deep learning techniques, including Graph Neural Networks (GNNs), have been employed to extract spatial relationships within traffic networks (Jin et al., 2023), while Transformer-based architectures have been utilized to capture temporal dependencies over various timescales (Shao et al., 2023a). Although these methods have shown promising results, they often rely on datasets that do not fully encapsulate the complexities introduced by rapid population growth and the surging number of vehicles, thus limiting their applicability to real-world scenarios.
There is an emerging need in intelligent transportation systems to design predictive models that extend beyond test adaptation, effectively generalizing to real-world conditions that evolve over time. It is important to note that our concept of ’beyond test adaptation’ differs from ’test time adaptation’ (Guo et al., 2024) as illustrated in Fig. 1 that shows the distinctions between them. This shift requires models that can handle the multifaceted impacts of demographic changes and vehicle proliferation without relying solely on adaptation to specific test datasets. To achieve this, it is essential to utilize datasets that accurately represent these evolving conditions over extremely long periods, capturing the intricate patterns influenced by population and vehicular growth.
Figure 1. Test-time adaptation in time-series forecasting involves training a single model to fit different test domains, horizons, or gaps. The figure above illustrates this using a gap example. In contrast, the figure below shows our ’beyond test adaptation’ where we train separate models for various gap settings.
Motivated by this need, we introduce XXLTraffic, a dataset and framework that expands traffic forecasting beyond test adaptation. By incorporating extremely long-term data, XXLTraffic better reflects real-world scenarios where traffic patterns are continually affected not just by infrastructure changes like highway construction, but also by shifts in distribution due to factors like population growth and increasing vehicle numbers. We will discuss existing datasets and the specific challenges encountered in establishing XXLTraffic, highlighting how it advances the field by providing a more realistic and comprehensive dataset. This facilitates the development of models capable of adapting to the complexities of real-world traffic dynamics without the limitations of traditional test adaptation approaches.
@@ 1. Introduction
Figure 2. Our dataset is evolving and longer than existing datasets. Existing datasets are either limited by short temporal spans or insufficient spatial nodes. In contrast, our dataset features an evolving growth of spatial nodes and spans over 20 years.
1.1. Recent Advances in Expanding Traffic Datasets
@@
Real-world traffic scenarios necessitate more complex prediction settings, involving extended temporal horizons or broader spatial coverage in experiments. In the temporal domain, new settings are typically proposed based on previously published work rather than introducing new datasets:  (Shao et al., 2023b) and (Jia et al., 2024) expanded input and output lengths up to four times on existing datasets.From a spatial perspective, (Chen et al., 2021) published a dataset with nodes growing annually and provided an evolving network to support new node predictions. (Wang et al., 2023b) proposed a continual learning framework with pattern expansion mechanisms based on (Chen et al., 2021). Additionally, SCPT (Prabowo et al., 2024) and Large-ST (Liu et al., 2024b) offered larger-scale spatial node datasets to support subsequent researchers. Recent work has explored longer temporal step experimental settings and released traffic datasets spanning up to five years and thousands of nodes. However, in specific scenarios, such as future traffic prediction for highway planning, these data and experimental settings fall short. As shown in Fig. 2, most existing datasets have limitations in temporal span, which inspired us to develop a dataset for expanding and extremely long traffic forecasting. This need for traffic forecasting beyond test adaptation is crucial in various real-world scenarios. For instance, urban planning and infrastructure investment decisions rely heavily on accurate long-term traffic predictions to ensure that developments meet future transportation demands. Commercial real estate site selection and development also depend on knowing future traffic volumes years in advance to optimize location choices and investment strategies. Additionally, governments can formulate more effective environmental policies based on long-term traffic forecasts, such as implementing traffic restrictions or promoting electric vehicles to reduce emissions. These applications highlight the importance of developing predictive models capable of accurately forecasting traffic patterns over extended periods, facilitating strategic decision-making across multiple sectors. As the temporal span extends, urban infrastructure development and road construction can lead to shifts in traffic patterns, resulting in an evolving domain shift. This observation motivated us to provide an expanding and extremely long traffic dataset. Additionally, the combination of these factors enables the extraction of more temporal patterns from extremely long sequences, allowing for the possibility of longer input sequences.
1.2. Challenges
@@
The ultra-dynamic challenge encompasses three key aspects: (1) Continuously evolving states of the underlying spatio-temporal infrastructures, characterized by an expanding number of nodes over the years. This continuous growth introduces complexity as the infrastructure adapts and expands. (2) Evolving temporal distributions over an extremely long observation period, which is crucial for extremely long forecasting beyond different noncontiguous train-test splits. This requires models to adapt to changes in patterns and trends over extensive temporal spans.
@@
We have constructed a traffic dataset with an exceptionally long temporal span and broader regional coverage, providing aggregated data and benchmarking, as well as a benchmarking setup considering extremely long prediction scenarios for future exploration:
@@
  • We present an experimental setup with temporal gaps for extremely long prediction beyong test adaptation and provide a benchmark of aggregated versions of hourly and daily datasets.
  • We provide exploration of input features through evolving temporal distributions over an extremely long observation period.
2. Preliminaries
@@
Traffic data primarily consists of vehicle flow detection data collected by sensors distributed across various locations in the traffic network. It is generally represented by $X_{i}$in $\mathbb{R}^{N\times T\times C}$, where $T$ denotes the time steps, $N$ denotes the number of sensors and $C$ denotes the number of features.
@@
(1) $[X_{t-(\alpha-1)},...,X_{t-1},X_{t}]\rightarrow[X_{t+1},X_{t+2},...,X_{t+\beta}],$
@@
This task mainly focuses on long-sequence time series prediction, which includes the traffic dataset. As shown in Table 1, the sequence length can reach up to 2880 steps.
@@
Based on Equation 1, the observation and prediction are not adjacent but instead are separated by a gap period $g$, as shown in the following formula.
(2) $[X_{t-(\alpha-1)},...,X_{t-1},X_{t}]\rightarrow[X_{t+g+1},X_{t+g+2},...,X_{t+g+\beta}],$
3. Gaps and Comparison with Existing Traffic Datasets
Figure 3. Our expanding and extremely long prediction addresses the existing limitations in both short-term and long-term predictions.
As shown in Table 1, existing traffic prediction work can easily be divided into short-term and long-term settings. The short-term setting originated from the STGCN(Yu et al., 2018) work, while the long-term setting was first introduced by LSTNet (Lai et al., 2018) and subsequently established as a widely adopted experimental framework by Informer (Zhou et al., 2021). In recent years, short-term prediction typically has a maximum step length of 12 steps, while long-term prediction reaches up to 720 steps. However, works such as Witran (Jia et al., 2024) and DAN (Li et al., 2024) recognized the need for even longer step predictions in practical applications, extending the length to a maximum of four times the typical length. Despite the differences in step lengths, their observed and predicted values are concatenated tightly together, as shown in Equation 1. To accommodate complex real-world scenarios, such as highway route planning predictions, it is necessary to introduce a gap of several years between observation and prediction. Typically, existing datasets lack the temporal coverage required to support gaps exceeding one year.At the same time, predicting several years in advance also implies the need to forecast traffic for sensors at new locations, taking into account the evolving nature of the road network. Even when such coverage is available, works like (Wang et al., 2023b) and (Chen et al., 2021) utilize evolving datasets but do not provide sufficient data to train models for extended durations. To overcome these gaps, our Expanding and Extremely Long Traffic Dataset robustly supports these complex scenarios.
Table 1. Summary of recent short-term traffic forecasting and long-term multivariate forecasting
@@ 3. Gaps and Comparison with Existing Traffic Datasets
Short-term STGCN (Yu et al., 2018) {3,6,9,12}
DCRNN (Li et al., 2018), GWN (Wu et al., 2019), BTF (Chen and Sun, 2021), {3,6,12}
DMSTGCN (Han et al., 2021), GTS (Shang et al., 2021), STGODE (Fang et al., 2021),
PM-MemNet (Lee et al., 2021), STAEFormer (Liu et al., 2023)
AGCRN (Bai et al., 2020), STSGCN (Song et al., 2020), ,DSTAGNN (Lan et al., 2022), {12}
D2STGNN (Shao et al., 2022), DyHSL (Zhao et al., 2023),PDFormer (Jiang et al., 2023),
MultiSPANS (Zou et al., 2024), GMSDR (Liu et al., 2022)
Long-term MTGNN (Wu et al., 2020a),LSTNet (Lai et al., 2018) {3,6,12,24}
ARU (Deshpande and Sarawagi, 2019) {12,24,48,168,336}
LogSparse_Trans (Li et al., 2019) {24,48,72,96,120,144,168,192}
AST (Wu et al., 2020b) {8,24,168,336}
SSDNet (Lin et al., 2021) {20,24,30,138}
Informer (Zhou et al., 2021), Autoformer (Wu et al., 2021), FEDformer (Zhou et al., 2022), {24,48,96,192,336,720}
Linear (Li et al., 2022), Triformer (Cirstea et al., 2022), Pyraformer (Liu et al., 2021)
DSformer (Yu et al., 2023),DeepTime (Woo et al., 2023),DLinear (Zeng et al., 2023)
Witran (Jia et al., 2024) {168, 336, 720, 1440, 2880}
DAN (Li et al., 2024) {288, 672, 1440}
4. The XXLTraffic Datasets
4.1. Data Collection
We obtained the expanding and extremely long traffic sensor data from the California Department of Transportation (CalTrans) Performance Measurement System [^1] https://pems.dot.ca.gov/ (PeMS) (Chen et al., 2001) and Transport for NSW [^2] https://maps.transport.nsw.gov.au/egeomaps/traffic-volumes/index.html. PeMS is an online platform that collects traffic data from 19,561 sensors distributed across California state highways. These sensor locations are divided into nine districts. We downloaded all the raw data for these nine districts from the initial data release up to March 20, 2024. The system automatically generates a daily data file for each district, containing data from all sensors within each district. We have stored the complete raw data files in an open-source repository for quick access, which will be released after the publication. The tfNSW is an open-source data platform provided by Transport for NSW, featuring traffic flow data collected from sensors distributed along major roads throughout the state of New South Wales of Australia. The data is available at a minimum granularity of one hour.
Table 2. Comparison between our XXLTraffic dataset and the existing traffic datasets.
@@ 4.1. Data Collection
DCRNN METR-LA 34,272 207 5 mins 4 months 03/2012 - 06/2012
@@
LSTNet Traffic 17,544 862 1 hour 2 years 01/2015 - 12/2016
STGCN PEMSD7(M) 12,672 228 5 mins 2 months 05/2012 - 06/2012
@@
ASTGCN PEMSD4-I 17,002 228 5 mins 2 months 01/2018 - 02/2018
@@
STSGCN PEMS03 26,208 358 5 mins 11 months 01/2018 - 11/2018
@@
Large-ST CA 525,888 8,600 5 mins 5 years 01/2017 - 12/2021
@@
tfNSW 100,056 27 60 mins 11.42 years 01/2013 - 05/2024
As illustrated in Table 2, our collected dataset significantly exceeds existing datasets in terms of both temporal coverage and the number of spatial nodes. The dataset sample will be available on: https://anonymous.4open.science/r/XXLTraffic-F281, which includes the raw data, sensor metadata (containing sensor IDs, geographical coordinates, associated road information, etc.), the data processing pipeline code, and the processed datasets.
4.2. Data Preprocessing
Based on the 23 years of raw data we collected, we conducted rigorous data filtering and aggregation. The PeMS system has continuously evolved, expanding from a few sensors in 2001 to over 4,000 sensors in some districts today. To support our setting of extremely long forecasting with gaps, we selected a subset of sensors that were installed in the early stages and have consistently collected new data up to the present(named gap dataset), which is shown in the Appendix. This extensive gap dataset effectively underpins the extremely long forecasting with gaps demonstrated in Fig. 3. Utilizing the gap dataset, we performed both hourly and daily aggregations, which will be employed for gap-free long-term forecasting benchmarking. We will provide standard long-term forecasting benchmarks for both the hourly and daily datasets.